### v1

# Sparse–Smooth Spatially Varying Coefficient Quantile Regression

Abstract
We develop a convex framework for spatially varying coefficient quantile regression that, for each predictor, separates a location-invariant global effect from a spatial deviation. An adaptive group penalty selects whether a predictor varies over space, while a graph–Laplacian quadratic promotes spatial continuity of the deviations on irregular networks. The formulation is identifiable via degree-weighted centering and scales with sparse linear algebra. We provide two practical solvers—an ADMM algorithm with closed-form proximal maps for the check loss and a smoothed proximal-gradient scheme based on the Moreau envelope—together with implementation guidance (projection for identifiability, stopping diagnostics, and preconditioning). Under mild conditions on the sampling design, covariates, error density, and graph geometry, we establish selection consistency for the deviation groups, mean-squared error bounds that balance Laplacian bias and stochastic variability, and root-$n$ asymptotic normality for the global coefficients with an oracle property. Simulations mimicking air-pollution applications demonstrate accurate recovery of global vs. local effects and competitive predictive performance under heteroskedastic, heavy-tailed noise. We discuss graph construction, spatially blocked cross-validation (to prevent leakage), and options for robust standard errors under spatial dependence.
keywords:
Quantile regression , spatially varying coefficients , graph Laplacian , group penalty , selection consistency , spatial cross-validation
journal: Spatial Statistics
[…]
organization=Center for Applied Statistics, School of Statistics, addressline=Renmin University of China, city=Beijing, postcode=100872, state=Beijing, country=China
[…]
organization=Center for Applied Statistics, School of Statistics, addressline=Renmin University of China, city=Beijing, postcode=100872, state=Beijing, country=China
[…]
organization=Center for Applied Statistics, School of Statistics, addressline=Renmin University of China, city=Beijing, postcode=100872, state=Beijing, country=China
{graphicalabstract}[Uncaptioned image]
{highlights}
Global–local decomposition for spatially varying coefficient quantile regression with adaptive $\ell_{2}$-group selection.
Graph-Laplacian smoothing on irregular networks with degree-weighted centering for identifiability.
Two scalable solvers: ADMM with closed-form proximal operators and a Moreau-smoothed proximal-gradient algorithm.
Selection consistency for spatial deviations and root-$n$ asymptotics for global coefficients (oracle property).
Practical guidance on graph construction, spatially blocked cross-validation, and robust standard errors.
1 Introduction
Spatial data often display nonstationary relations between a response and its predictors, so that regression coefficients vary over space. The geographically weighted regression (GWR) literature formalizes this idea via locally weighted fits across locations [Fotheringham2002]. Extensions include multiscale GWR to allow distinct spatial scales across coefficients [FotheringhamYangKang2017, OshanFotheringham2019]. In parallel, quantile regression (QR) [KoenkerBassett1978, Koenker2005] models conditional quantiles and is known to be robust under heavy tails and heteroskedasticity. Combining spatially varying coefficients (SVCs) with QR yields a tool for tail behavior and distributional heterogeneity. Bayesian and frequentist SVC–QR formulations have been studied using Gaussian-process priors or nonparametric components [Gelfand2003SVC, Reich2012STQR], but scalable convex formulations that simultaneously deliver interpretable global-vs-local decisions and graph-based smoothness remain relatively limited.
We propose a penalized SVC–QR estimator with two desiderata. First, for each covariate we separate a global component (constant in space) from a spatial deviation, and we use a group $L_{2}$ penalty to shrink the deviation to zero whenever the covariate is effectively global. This provides an interpretable global-vs-local decision not available in kernel-weighted local QR. Second, conditioned on being local, we regularize the deviation by a graph Laplacian quadratic constructed from the sampling locations. This encourages spatial continuity and can accommodate relatively sharp local changes when supported by the data; strictly edge-preserving behavior would require total-variation-type penalties, which we discuss as an extension. The formulation is convex and compatible with large-scale QR solvers, including proximal-gradient implementations under a smooth surrogate and ADMM schemes tailored to the check loss.
On theory, we show that global coefficients are root-$n$ consistent and asymptotically normal, deviation fields converge in mean square at a rate balancing Laplacian bias and stochastic variability, and selection of globally constant effects is consistent under a groupwise beta-min condition. Our assumptions allow random sampling locations filling a compact set, bounded covariates, and sparse graphs with uniformly bounded average degree or controlled growth.
We emphasize choices for reproducibility: mutual $k$-nearest-neighbor or compact-kernel graphs, spatially blocked cross-validation (to mitigate leakage), and diagnostics for KKT conditions and stopping. We also document computational scaling and preconditioning for large geospatial datasets.
2 Model
We develop a sparse–smooth spatially varying coefficient quantile regression (SS–SVCQR) at a fixed quantile $\tau\in(0,1)$. Scalars are in italic lower case, vectors in bold lower case, and matrices in bold upper case. For $\bm{v}$, $\|\bm{v}\|_{2}$ is the Euclidean norm, and $\bm{1}$ is an all-ones vector of conformable length.
We observe $\{(Y_{i},\bm{Z}_{i},\bm{X}_{i},u_{i}):i=1,\dots,n\}$, with $Y_{i}\in\mathbb{R}$, $\bm{Z}_{i}\in\mathbb{R}^{q}$ (purely global effects), $\bm{X}_{i}\in\mathbb{R}^{p}$ (potentially varying over space), and $u_{i}\in\mathcal{U}\subset\mathbb{R}^{2}$ the location. Let $Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})$ denote the conditional $\tau$-quantile. The target relation is
$Q_{\tau}\{Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i}\}=\bm{Z}_{i}^{\top}\bm{\alpha}+\sum_{j=1}^{p}X_{ij}\beta_{j}(u_{i}),\qquad i=1,\dots,n,$ (1)
with $\bm{\alpha}\in\mathbb{R}^{q}$ and spatially varying coefficients $\beta_{j}(\cdot)$. Each coefficient is decomposed into a global level and a spatial deviation,
$\beta_{j}(u)=\beta_{G,j}+\delta_{j}(u),\qquad j=1,\dots,p,$ (2)
where $\bm{\beta}_{G}=(\beta_{G,1},\dots,\beta_{G,p})^{\top}\in\mathbb{R}^{p}$ and $\delta_{j}:\mathcal{U}\to\mathbb{R}$.
The model is anchored at level $\tau$ by the quantile exogeneity condition:
$\Pr\big\{Y_{i}\leq\bm{Z}_{i}^{\top}\bm{\alpha}_{0}+\sum_{j=1}^{p}X_{ij}\beta_{0,j}(u_{i})\ \big|\ \bm{Z}_{i},\bm{X}_{i},u_{i}\big\}=\tau\quad\text{almost surely for each }i.$ (3)
Here $\bm{\beta}_{0}=\{\beta_{0,1}(\cdot),\dots,\beta_{0,p}(\cdot)\}$ is represented by $(\bm{\beta}_{G0},\{\delta_{0,j}(\cdot)\}_{j=1}^{p})$ with $\beta_{0,j}(u)=\beta_{G0,j}+\delta_{0,j}(u)$.
Spatial geometry is encoded by a symmetric graph on $\{u_{i}\}$. Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric adjacency from a mutual $k$-nearest-neighbor rule or a symmetric compact-support kernel, and let $\bm{D}=\operatorname{diag}(\bm{A}\bm{1})$ collect node degrees. We use the symmetric normalized Laplacian
$\bm{L}_{\mathrm{sym}}=\bm{I}-\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2},$ (4)
which is symmetric positive semidefinite with spectrum in $[0,2]$.
To make (2) unique and prevent drift between global and local parts, local deviations are degree-weight centered componentwise. For a component $\mathcal{C}\subset\{1,\dots,n\}$, let $\bm{1}_{\mathcal{C}}$ be its indicator. Writing $\bm{\delta}_{j}=(\delta_{j}(u_{1}),\dots,\delta_{j}(u_{n}))^{\top}$, the identifiability constraint is
$\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ (5)
for every connected component $\mathcal{C}$ and all $j$. Because $\mathrm{null}(\bm{L}_{\mathrm{sym}})$ on a component is spanned by $\bm{D}^{1/2}\bm{1}_{\mathcal{C}}$, orthogonality of $\bm{D}^{1/2}\bm{\delta}_{j}$ to the null space is equivalent to (5). Under (5), $\beta_{G,j}$ equals the degree-weighted average of the spatial coefficient $\beta_{j}(\cdot)$ across nodes; $\delta_{j}(\cdot)$ captures pure spatial departures around that average.
Vectorization at observed sites fixes the linear operators. Stack $\bm{y}=(Y_{1},\dots,Y_{n})^{\top}$, $\bm{Z}=[\bm{Z}_{1}^{\top};\cdots;\bm{Z}_{n}^{\top}]\in\mathbb{R}^{n\times q}$, and $\bm{X}=[\bm{X}_{1}^{\top};\dots;\bm{X}_{n}^{\top}]\in\mathbb{R}^{n\times p}$. For $j=1,\dots,p$, define $\bm{X}_{\odot j}=\operatorname{diag}(X_{1j},\dots,X_{nj})\in\mathbb{R}^{n\times n}$ and $\bm{\delta}_{j}\in\mathbb{R}^{n}$. Concatenate deviations by
$\bm{\Delta}=\big[\bm{\delta}_{1}^{\top},\ \bm{\delta}_{2}^{\top},\ \dots,\ \bm{\delta}_{p}^{\top}\big]^{\top}\in\mathbb{R}^{np},$
and build the block operator
$\bm{X}_{\delta}=\big[\bm{X}_{\odot 1}\ \bm{X}_{\odot 2}\ \cdots\ \bm{X}_{\odot p}\big]\in\mathbb{R}^{n\times np}.$ (6)
The conditional quantile vector implied by (1)–(2) is
$\bm{q}_{\tau}=\bm{Z}\bm{\alpha}+\bm{X}\bm{\beta}_{G}+\bm{X}_{\delta}\bm{\Delta},$ (7)
and the residuals at $\theta=(\bm{\alpha},\bm{\beta}_{G},\bm{\Delta})$ are $\bm{r}(\theta)=\bm{y}-\bm{q}_{\tau}$. The local component $\bm{X}_{\delta}\bm{\Delta}=\sum_{j=1}^{p}\bm{X}_{\odot j}\bm{\delta}_{j}$ is a superposition of $p$ diagonal-scaled fields.
Spatial smoothness is summarized by
$\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j}=\tfrac{1}{2}\sum_{(i,\ell)\in\mathcal{E}}A_{i\ell}\Big\{\tfrac{\delta_{j}(u_{i})}{\sqrt{d_{i}}}-\tfrac{\delta_{j}(u_{\ell})}{\sqrt{d_{\ell}}}\Big\}^{2},\qquad d_{i}=(\bm{D})_{ii},$ (8)
where $\mathcal{E}$ is the set of undirected edges (each unordered pair counted once). The quantity vanishes only for degree-weighted constants on each component.
Interpretation follows: a predictor $X_{:j}$ is purely global when $\bm{\delta}_{j}=\bm{0}$ and $\beta_{G,j}\neq 0$, purely local with zero degree-weighted mean when $\beta_{G,j}=0$ and $\bm{\delta}_{j}\neq\bm{0}$, and mixed when both parts are nonzero. Disconnected graphs are handled componentwise without changing notation.
3 Estimation and Algorithms
We estimate $(\bm{\theta}_{P},\bm{\Theta}_{L})$ with $\bm{\theta}_{P}=(\bm{\alpha}^{\top},\bm{\beta}_{G}^{\top})^{\top}\in\mathbb{R}^{q+p}$ and $\bm{\Theta}_{L}=\{\bm{\delta}_{j}\in\mathbb{R}^{n}:j=1,\dots,p\}$, each $\bm{\delta}_{j}$ degree-weight centered on every component.
The empirical objective at $\tau$ is
$\min_{\bm{\theta}_{P},\bm{\Theta}_{L}}\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$ (9)
subject to $\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ for every component $\mathcal{C}$ and all $j$. Weights $w_{j}>0$ enable adaptive grouping, e.g. $w_{j}=(\|\tilde{\bm{\delta}}_{j}\|_{2}+a)^{-\gamma}$ with $a>0$ and $\gamma\in(0,1]$. Tuning $(\lambda_{1},\lambda_{2})$ is selected by spatially blocked cross-validation.
Define $\bm{r}(\bm{\theta}_{P},\bm{\Theta}_{L})=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\bm{X}_{\delta}\bm{\Delta}$ with $\bm{\Delta}=[\bm{\delta}_{1}^{\top},\dots,\bm{\delta}_{p}^{\top}]^{\top}$. The objective (9) is convex and separable across groups for the non-smooth term.
Handling the centering constraint.
After each update of $\bm{\delta}_{j}$, project it to the degree-weight centered subspace. Let $\{\mathcal{C}_{k}\}_{k=1}^{K}$ be the components and $\bm{1}_{\mathcal{C}_{k}}$ their indicators. The projector reads
$\mathsf{Proj}_{D}(\bm{v})=\bm{v}-\sum_{k=1}^{K}\frac{\bm{1}_{\mathcal{C}_{k}}^{\top}\bm{D}\bm{v}}{\bm{1}_{\mathcal{C}_{k}}^{\top}\bm{D}\bm{1}_{\mathcal{C}_{k}}}\bm{1}_{\mathcal{C}_{k}},$ (10)
and we enforce $\bm{\delta}_{j}\leftarrow\mathsf{Proj}_{D}(\bm{\delta}_{j})$ every iteration.
An ADMM scheme with closed-form proximal maps
Introduce $\bm{s}=\bm{r}(\bm{\theta}_{P},\bm{\Theta}_{L})$ and $\bm{z}_{j}=\bm{\delta}_{j}$ for $j=1,\dots,p$, and solve
$\min_{\bm{\theta}_{P},\bm{\Theta}_{L},\bm{s},\{\bm{z}_{j}\}}\sum_{i=1}^{n}\rho_{\tau}(s_{i})+\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{z}_{j}\|_{2}+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j}$ (11)
subject to $\bm{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j=1}^{p}\bm{X}_{\odot j}\bm{\delta}_{j}$ and $\bm{z}_{j}=\bm{\delta}_{j}$, together with centering via (10). Using scaled duals $(\bm{u},\{\bm{v}_{j}\})$ and penalties $(\rho_{s},\rho_{z})>0$, one ADMM sweep is:
Update of $(\bm{\alpha},\bm{\beta}_{G})$. Solve
$\left[\begin{array}[]{cc}\bm{Z}^{\top}\bm{Z}&\bm{Z}^{\top}\bm{X}\\ \bm{X}^{\top}\bm{Z}&\bm{X}^{\top}\bm{X}\end{array}\right]\left[\begin{array}[]{c}\bm{\alpha}\\ \bm{\beta}_{G}\end{array}\right]=\left[\begin{array}[]{c}\bm{Z}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\\[2.0pt] \bm{X}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\end{array}\right].$
The coefficient matrix depends only on $\bm{Z}\ \bm{X}]$ and can be factorized once.
Update of $\bm{s}$. The proximal map of the check loss is an asymmetric soft-threshold. For each $i$,
$\bm{s}\leftarrow\mathrm{prox}_{\rho_{s}^{-1}\rho_{\tau}}(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}+\bm{u}),$ (12)
with
$\mathrm{prox}_{\gamma\rho_{\tau}}(v)=\begin{cases}v-\gamma\tau,&v>\gamma\tau,\\ 0,&-\gamma(1-\tau)\leq v\leq\gamma\tau,\\ v+\gamma(1-\tau),&v<-\gamma(1-\tau).\end{cases}$
Update of $\bm{\delta}_{j}$. Holding other blocks fixed, $\bm{\delta}_{j}$ solves
$\big(2\lambda_{2}\bm{L}_{\mathrm{sym}}+\rho_{s}\bm{X}_{\odot j}^{\top}\bm{X}_{\odot j}+\rho_{z}\bm{I}\big)\bm{\delta}_{j}=\rho_{s}\bm{X}_{\odot j}^{\top}\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{\ell\neq j}\bm{X}_{\odot\ell}\bm{\delta}_{\ell}-\bm{s}+\bm{u}\big)+\rho_{z}(\bm{z}_{j}-\bm{v}_{j}).$ (13)
Solve by sparse Cholesky or conjugate gradients; then project $\bm{\delta}_{j}\leftarrow\mathsf{Proj}_{D}(\bm{\delta}_{j})$.
Update of $\bm{z}_{j}$. Apply group soft-thresholding,
$\bm{z}_{j}\leftarrow\Big(1-\frac{\lambda_{1}w_{j}}{\rho_{z}\|\bm{\delta}_{j}+\bm{v}_{j}\|_{2}}\Big)_{+}(\bm{\delta}_{j}+\bm{v}_{j}),\qquad j=1,\dots,p.$ (14)
Dual updates and stopping. Update
$\bm{u}\leftarrow\bm{u}+\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}\big),\qquad\bm{v}_{j}\leftarrow\bm{v}_{j}+(\bm{\delta}_{j}-\bm{z}_{j}),$
and stop when primal residuals
$\bm{r}_{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s},\qquad\bm{r}_{z,j}=\bm{\delta}_{j}-\bm{z}_{j}$
and dual residuals
$\bm{d}_{s}=\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}}),\qquad\bm{d}_{z,j}=\rho_{z}(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}})$
satisfy $\|\bm{r}_{s}\|_{2}\leq\varepsilon_{\mathrm{pri}}$, $\big(\sum_{j}\|\bm{r}_{z,j}\|_{2}^{2}\big)^{1/2}\leq\varepsilon_{\mathrm{pri}}$, $\|\bm{d}_{s}\|_{2}\leq\varepsilon_{\mathrm{dual}}$ and $\big(\sum_{j}\|\bm{d}_{z,j}\|_{2}^{2}\big)^{1/2}\leq\varepsilon_{\mathrm{dual}}$, where tolerances scale with $\sqrt{n}$ and block magnitudes. Over-relaxation and adaptive penalty updates are often beneficial.
A smoothed proximal–gradient scheme with Moreau smoothing
Define the smoothed loss $M_{h}(r)=\min_{s}\{\rho_{\tau}(s)+\tfrac{1}{2h}(s-r)^{2}\}$. Its gradient exists and is $1/h$-Lipschitz:
$\nabla_{r}M_{h}(r)=\frac{1}{h}\big(r-\mathrm{prox}_{h\rho_{\tau}}(r)\big).$
Let
$\mathcal{G}_{h}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}M_{h}\Big(r_{i}(\bm{\theta}_{P},\bm{\Theta}_{L})\Big)+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},\qquad\mathcal{R}(\bm{\Theta}_{L})=\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}.$
Gradients follow by the chain rule:
$\nabla_{\bm{\alpha}}\mathcal{G}_{h}=-\bm{Z}^{\top}\bm{g},\quad\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h}=-\bm{X}^{\top}\bm{g},\quad\nabla_{\bm{\delta}_{j}}\mathcal{G}_{h}=-\bm{X}_{\odot j}^{\top}\bm{g}+2\lambda_{2}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
with $\bm{g}=\frac{1}{h}(\bm{r}-\mathrm{prox}_{h\rho_{\tau}}(\bm{r}))$. Given a step size $t>0$ (via backtracking to ensure sufficient decrease), update
$\bm{\alpha}\leftarrow\bm{\alpha}-t\nabla_{\bm{\alpha}}\mathcal{G}_{h},\qquad\bm{\beta}_{G}\leftarrow\bm{\beta}_{G}-t\nabla_{\bm{\beta}_{G}}\mathcal{G}_{h},$
$\bm{\delta}_{j}\leftarrow\operatorname{shrink}_{t\lambda_{1}w_{j}}\Big(\bm{\delta}_{j}-t\nabla_{\bm{\delta}_{j}}\mathcal{G}_{h}\Big),\quad\operatorname{shrink}_{\kappa}(\bm{v})=\Big(1-\frac{\kappa}{\|\bm{v}\|_{2}}\Big)_{+}\bm{v},$
followed by $\bm{\delta}_{j}\leftarrow\mathsf{Proj}_{D}(\bm{\delta}_{j})$. Nesterov acceleration with standard restart is used. A continuation $h\downarrow h_{\min}$ with $\sqrt{n}h\to\infty$ preserves asymptotics.
Tuning, initialization, and implementation details
Tuning $(\lambda_{1},\lambda_{2})$ and graph hyperparameters (e.g. mutual $k$ in $k$-NN or kernel bandwidth) uses spatially blocked cross-validation. Partition locations into $K$ spatial folds; for each held-out fold, construct the graph and Laplacian using only the training folds, fit the model, and evaluate the held-out check loss. This prevents leakage through cross-fold edges. Adaptive weights use a pilot fit with small $\lambda_{1}$ and moderate $\lambda_{2}$, or a purely smooth fit with $\lambda_{1}=0$; set $w_{j}=(\|\tilde{\bm{\delta}}_{j}\|_{2}+a)^{-\gamma}$ with a small $a$ for stability. Initialize $(\bm{\alpha},\bm{\beta}_{G})$ at the global QR and set $\bm{\delta}_{j}=\bm{0}$.
Both schemes exploit sparsity. $\bm{L}_{\mathrm{sym}}$ is stored in a compressed sparse format; matrix–vector products cost $O(nd)$ with average degree $d$. The multipliers $\bm{X}_{\odot j}$ apply in $O(n)$. In ADMM, systems in (13) have the form $\rho_{z}\bm{I}+\rho_{s}\mathrm{diag}(X_{:j}^{2})+2\lambda_{2}\bm{L}_{\mathrm{sym}}$ and admit efficient preconditioned CG with a shared preconditioner (e.g. incomplete Cholesky of $\rho_{z}\bm{I}+2\lambda_{2}\bm{L}_{\mathrm{sym}}$). In the smoothed scheme, costs are dominated by proximal evaluations (linear time) and sparse graph multiplications.
Stopping rules and KKT diagnostics
For ADMM, stopping relies on primal/dual residuals as above, with tolerances
$\varepsilon_{\mathrm{pri}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\{\|\bm{s}\|_{2},\ \|\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}\|_{2}\},$
$\varepsilon_{\mathrm{dual}}=\sqrt{n}\varepsilon_{\mathrm{abs}}+\varepsilon_{\mathrm{rel}}\max\{\|\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}})\|_{2},\ (\sum_{j}\|\rho_{z}(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}})\|_{2}^{2})^{1/2}\}.$
For the smoothed scheme, stopping uses relative objective decrease and a KKT residual,
$\big\|\bm{Z}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\big\|_{2}+\big\|\bm{X}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\big\|_{2}+\sum_{j=1}^{p}\mathrm{dist}\Big(\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})+2\lambda_{2}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j},\ \lambda_{1}w_{j}\partial\|\hat{\bm{\delta}}_{j}\|_{2}\Big),$
where $\bm{\psi}_{\tau}(\bm{r})$ stacks $\psi_{\tau}(r_{i})=\tau-\mathbf{1}\{r_{i}<0\}$. The distance to the group subdifferential is
$\mathrm{dist}(\bm{v},\ \lambda_{1}w_{j}\partial\|\hat{\bm{\delta}}_{j}\|_{2})=\begin{cases}\left\|\bm{v}-\lambda_{1}w_{j}\dfrac{\hat{\bm{\delta}}_{j}}{\|\hat{\bm{\delta}}_{j}\|_{2}}\right\|_{2},&\hat{\bm{\delta}}_{j}\neq 0,\\[6.0pt] \max\{\|\bm{v}\|_{2}-\lambda_{1}w_{j},\ 0\},&\hat{\bm{\delta}}_{j}=0.\end{cases}$
Practical choices
Penalty scales can be anchored by heuristics. A convenient grid sets $\lambda_{2}$ around the median of the nonzero eigenvalues of $\bm{L}_{\mathrm{sym}}$ divided by $n$, and $\lambda_{1}$ around the empirical $(1-\tau)\tau$ score scale times $\sqrt{\log p/n}$. In noisy data, a continuation schedule that starts with larger $\lambda_{2}$ and decreases gradually stabilizes both solvers. When repeated locations occur, connect coincident points by a high-weight edge or merge them.
4 Asymptotic Theory
We analyze the SS–SVCQR estimator in Section 2. The parametric block is $\bm{\theta}_{P}=(\bm{\alpha}^{\top},\bm{\beta}_{G}^{\top})^{\top}\in\mathbb{R}^{q+p}$, and the local deviations are $\bm{\Theta}_{L}=\{\bm{\delta}_{j}\in\mathbb{R}^{n}:j=1,\dots,p\}$, each degree-weight centered on every graph component.
Estimator and penalty.
At quantile level $\tau$,
$\mathcal{L}_{n}(\bm{\theta}_{P},\bm{\Theta}_{L})=\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\sum_{j=1}^{p}\lambda_{1n}w_{j,n}\|\bm{\delta}_{j}\|_{2}+\sum_{j=1}^{p}\lambda_{2n}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
subject to $\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ for every component $\mathcal{C}$ and every $j$. Weights $w_{j,n}>0$ can be adaptive, e.g. $w_{j,n}=(\|\tilde{\bm{\delta}}_{j}\|_{2}+a_{n})^{-\gamma}$ with $a_{n}\downarrow 0$, $\gamma\in(0,1]$.
Population target and notation.
Let $(\bm{\alpha}_{0},\bm{\beta}_{0})$ denote the true parameters in the sense of (3). Write $\beta_{0,j}(u)=\beta_{G0,j}+\delta_{0,j}(u)$ with degree-weight centered $\delta_{0,j}(\cdot)$ and define
$\bm{\delta}_{j0}=\big(\delta_{0,j}(u_{1}),\dots,\delta_{0,j}(u_{n})\big)^{\top},\qquad\mathcal{A}_{0}=\{j:\bm{\delta}_{j0}=\bm{0}\},\ \ \mathcal{A}_{0}^{c}=\{1,\dots,p\}\setminus\mathcal{A}_{0}.$
Assumption 1 (Sampling design and graph).
Locations $u_{i}$ are i.i.d. with a density bounded away from zero and infinity on a compact $\mathcal{U}\subset\mathbb{R}^{2}$. The graph is constructed either (i) by a symmetric mutual $k_{n}$-nearest-neighbor rule with $k_{n}$ bounded away from zero and $k_{n}=O(1)$ (node-level estimation focus), or (ii) by a symmetric compact-support kernel with bandwidth bounded away from zero. With probability tending to one the graph has $O(1)$ connected components. Degrees satisfy $\max_{i}d_{i}=O_{p}(1)$ and $\min_{i}d_{i}\geq c>0$ with high probability.
Remark 1.
If one targets continuum limits of $\delta_{j}(\cdot)$, a growing-degree regime can be imposed (e.g., $k_{n}\to\infty$, $k_{n}/n\to 0$ or shrinking bandwidth), together with a suitable scaling of the Laplacian; our theory focuses on node-level estimation and keeps $d_{i}=O_{p}(1)$.
Assumption 2 (Covariates and Gram matrices).
Covariates are uniformly bounded almost surely, $\|\bm{Z}_{i}\|_{\infty}\leq C_{Z}$ and $\|\bm{X}_{i}\|_{\infty}\leq C_{X}$. The Gram matrices $n^{-1}\bm{Z}^{\top}\bm{Z}$ and $n^{-1}\bm{X}^{\top}\bm{X}$ converge to positive definite limits. For each $j$, $\bm{X}_{\odot j}$ is not identically zero on any component.
Assumption 3 (Errors and local density).
Let $\varepsilon_{i}=Y_{i}-Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})$. Then $\Pr(\varepsilon_{i}\leq 0\mid\bm{Z}_{i},\bm{X}_{i},u_{i})=\tau$ a.s., the conditional density $f_{\varepsilon\mid Z,X,U}(t)$ exists and is continuous near $t=0$, and $0<c_{f}\leq f_{\varepsilon\mid Z,X,U}(0)\leq C_{f}<\infty$ a.s.
Assumption 4 (Smoothness of local deviations).
For each $j\in\mathcal{A}_{0}^{c}$, the true deviation vector satisfies $\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\leq C_{j}$ with constants $C_{j}<\infty$ not depending on $n$.
Assumption 5 (Penalty sequences and adaptive weights).
The smoothness penalty obeys $\lambda_{2n}\to 0$ and $n\lambda_{2n}\to\infty$. The sparsity penalty satisfies $\lambda_{1n}\to 0$ and $\sqrt{n}\lambda_{1n}\to\infty$. Adaptive weights obey $\sup_{j\in\mathcal{A}_{0}^{c}}w_{j,n}=O_{p}(1)$ and $\inf_{j\in\mathcal{A}_{0}}w_{j,n}\geq c_{w}>0$ with probability tending to one.
Assumption 6 (Beta–min separation).
There exists $c_{\beta}>0$ such that for all $j\in\mathcal{A}_{0}^{c}$, $\|\bm{\delta}_{j0}\|_{2}\geq c_{\beta}\lambda_{1n}w_{j,n}$ for all sufficiently large $n$ with probability tending to one.
Assumption 7 (Restricted curvature for the quantile loss).
Let $\bm{G}_{i}=(\bm{Z}_{i}^{\top},\bm{X}_{i}^{\top})^{\top}$ and define
$\bm{M}=\mathrm{E}\big[f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}\big],\qquad\bm{V}=\tau(1-\tau)\mathrm{E}\big[\bm{G}_{i}\bm{G}_{i}^{\top}\big].$
Then $\bm{M}$ is positive definite. Moreover, for any $\bm{v}$ in a neighborhood of the origin,
$n^{-1}\sum_{i=1}^{n}\big\{\rho_{\tau}(\varepsilon_{i}-\bm{G}_{i}^{\top}\bm{v})-\rho_{\tau}(\varepsilon_{i})\big\}=\frac{1}{2}\bm{v}^{\top}\bm{M}\bm{v}+o_{p}(\|\bm{v}\|_{2}^{2}),$
uniformly over events where $\|\bm{v}\|_{2}=O(n^{-1/2})$.
Remark 2.
Our asymptotics cover fixed $p$. For growing $p$, additional group-sparsity conditions (e.g., restricted eigenvalues) are required and are left for future work.
KKT characterization.
Let $\psi_{\tau}(r)=\tau-\mathbf{1}\{r<0\}$ and $\hat{r}_{i}$ be the fitted residuals. At any global minimizer $(\hat{\bm{\theta}}_{P},\hat{\bm{\Theta}}_{L})$, the parametric block satisfies
$\sum_{i=1}^{n}\psi_{\tau}(\hat{r}_{i})\bm{G}_{i}=\bm{0}.$
For each $\bm{\delta}_{j}$,
$\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})+2\lambda_{2n}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}+\bm{D}\bm{\mu}_{j}\in-\lambda_{1n}w_{j,n}\partial\|\hat{\bm{\delta}}_{j}\|_{2},$
with centering constraints $\bm{1}_{\mathcal{C}}^{\top}\bm{D}\hat{\bm{\delta}}_{j}=0$ on every component and multipliers $\bm{\mu}_{j}$.
Deviation rates and selection consistency
Theorem 3 (Deviation rates and selection consistency).
Suppose Assumptions 16 hold. Then:
  1. 1.
    For each $j\in\mathcal{A}_{0}^{c}$,
    $\frac{1}{n}\sum_{i=1}^{n}\big\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\big\}^{2}=O_{p}\Big(\frac{1}{n\lambda_{2n}}+\lambda_{2n}\mathcal{S}_{j}\Big),$
    where $\mathcal{S}_{j}$ depends on $\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}$ and the spectrum of $\bm{L}_{\mathrm{sym}}$. The bound is balanced at $\lambda_{2n}\asymp(n\mathcal{S}_{j})^{-1/2}$.
  2. 2.
    The adaptive group penalty correctly separates globally constant and truly varying covariates:
    $\Pr\big(\hat{\bm{\delta}}_{j}=\bm{0}\ \text{for all }j\in\mathcal{A}_{0}\big)\to 1,\qquad\Pr\big(\hat{\bm{\delta}}_{j}\neq\bm{0}\ \text{for all }j\in\mathcal{A}_{0}^{c}\big)\to 1.$
    On the event of correct selection, the deviations satisfy the rate in part (i).
Proof sketch. A localized quadratic expansion of the empirical check loss in the direction of $\bm{\delta}_{j}$, combined with the Laplacian quadratic, yields a linear approximation whose stochastic fluctuation in the degree-weight centered subspace is of order $(n\lambda_{2n})^{-1/2}$. The deterministic bias is of order $\lambda_{2n}\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}$. Selection follows by a primal–dual witness argument together with the bound $\|\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{r}_{0})\|_{2}=O_{p}(\sqrt{n})$ under bounded covariates.
Parametric limit distribution and oracle property
Theorem 4 (Asymptotic normality and oracle property).
Suppose Assumptions 17 hold. Then
$\sqrt{n}(\hat{\bm{\theta}}_{P}-\bm{\theta}_{P0})\ \overset{d}{\longrightarrow}\ \mathcal{N}\big(\bm{0},\ \bm{M}^{-1}\bm{V}\bm{M}^{-1}\big).$
Moreover, on $\{\hat{\mathcal{A}}=\mathcal{A}_{0}\}$ (which holds with probability tending to one by Theorem 3), the joint limit of $(\hat{\bm{\theta}}_{P},\{\hat{\bm{\delta}}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ coincides with that of the oracle estimator that knows $\bm{\delta}_{j0}=\bm{0}$ for all $j\in\mathcal{A}_{0}$.
Lemma 1 (Cross-term control).
Under bounded covariates and fixed $p$, if $\lambda_{2n}\asymp n^{-1/2}$, then
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}(\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i}))\Big)\bm{G}_{i}\right\|_{2}=o_{p}(1).$
Sketch. By Cauchy–Schwarz and boundedness,
$\lesssim\sqrt{\frac{1}{n}\sum_{i=1}^{n}\Big(\sum_{j}X_{ij}e_{j,i}\Big)^{2}}\ \lesssim\ \sqrt{\frac{1}{n}\sum_{i,j}e_{j,i}^{2}}\ =\ O_{p}\left(\sqrt{\frac{1}{n}\sum_{j}\mathrm{MSE}_{\delta,j}}\right)=\ O_{p}(n^{-1/4})\to 0,$
using Theorem 3(i). This yields a negligible remainder in the Bahadur expansion of $\hat{\bm{\theta}}_{P}$.
Standard error estimation.
Estimate the asymptotic variance using the plug-in
$\hat{\bm{M}}=n^{-1}\sum_{i=1}^{n}\hat{f}_{i}(0)\bm{G}_{i}\bm{G}_{i}^{\top},\qquad\hat{\bm{V}}=\tau(1-\tau)n^{-1}\sum_{i=1}^{n}\bm{G}_{i}\bm{G}_{i}^{\top},$
where $\hat{f}_{i}(0)$ is a local estimator of $f_{\varepsilon\mid Z,X,U}(0)$ (e.g., kernel density on residuals) or via a smoothed objective with bandwidth $h_{n}\downarrow 0$ and $\sqrt{n}h_{n}\to\infty$. When spatial dependence is a concern, one may use spatial block bootstrap or Conley-type HAC corrections in the sandwich step.
5 Simulation Study
We conducted two complementary simulation experiments. Scenario 1 uses a single relatively dense data set to provide a visual check that the proposed estimator can correctly recover the spatial structure of the coefficients under a variety of error distributions. Scenario 2 is a Monte Carlo study that summarizes selection accuracy, estimation error, and predictive performance over repeated samples.
Both scenarios share the same basic design for spatial locations, covariates, and true coefficient surfaces; only the sample size, error distribution, and the way the results are summarized differ. The estimation algorithm (ADMM implementation of SS–SVCQR with two-stage adaptive group penalty and spatially blocked cross-validation) is exactly as described in Section 3 and is therefore not repeated here.
Settings
We place $n$ monitoring sites on the unit square $\mathcal{U}=[0,1]^{2}$. The locations $u_{i}=(u_{i1},u_{i2})^{\top}$ are i.i.d. $\mathrm{Unif}([0,1]^{2})$, mimicking an irregular but dense network of stations. At each site, we observe a global block $\bm{Z}_{i}\in\mathbb{R}^{3}$ and a potentially spatially varying block $\bm{X}_{i}\in\mathbb{R}^{4}$,
$Z_{i0}\equiv 1,\quad Z_{i1}\sim\mathcal{N}(0,1),\quad Z_{i2}\sim\mathcal{N}(0,1),\quad X_{ij}\sim\mathcal{N}(0,1)\ (j=1,\dots,4),$
independently in $i$ and $j$.
[…]
$Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})=\bm{Z}_{i}^{\top}\bm{\alpha}_{0}+\sum_{j=1}^{4}X_{ij}\beta_{0,j}(u_{i}),\quad\beta_{0,j}(u)=\beta_{G0,j}+\delta_{0,j}(u),$
with global effects
$\bm{\alpha}_{0}=(3,-1.0,1.5)^{\top},\qquad\bm{\beta}_{G0}=(5,0,2.5,0)^{\top},$
so that $X_{1}$ and $X_{3}$ have nonzero global components, while $X_{2}$ and $X_{4}$ would be purely global if their spatial deviations vanished.
We design the deviation fields so that exactly two covariates are truly local:
$\delta_{0,1}(u)$ $=A_{1}\left\{\sin(2\pi u_{1})\cos(2\pi u_{2})+c_{1}(u_{1}-0.5)\right\},$
$\delta_{0,2}(u)$ $\equiv 0,$
$\delta_{0,3}(u)$ $=A_{3}\left\{(u_{1}-0.5)^{2}+(u_{2}-0.5)^{2}\right\},$
$\delta_{0,4}(u)$ $\equiv 0,$
where $(A_{1},c_{1},A_{3})$ are fixed constants chosen so that the signal amplitude is comparable to the random noise. In the implementation we evaluate these fields at the observed sites, construct a $k$-nearest-neighbour graph on $\{u_{i}\}$, and apply the same degree-weighted centering operator as in Section 2 so that each $\delta_{0,j}(\cdot)$ has zero degree-weighted mean on every connected component. Under this centering, $\beta_{G0,j}$ is the degree-weighted spatial average of $\beta_{0,j}(\cdot)$ and $\delta_{0,j}(\cdot)$ represents pure spatial departures.
5.1 Recovery of Spatial Structure
Scenario 1 uses a single, relatively large dataset ($n=1000$) for each error distribution and focuses on qualitative recovery of the spatial deviation fields. Responses are generated as
$Y_{i}=Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})+\varepsilon_{i},$
with $\varepsilon_{i}$ independent of $(\bm{Z}_{i},\bm{X}_{i},u_{i})$.
We consider several error distributions to probe robustness:
  1. 1.
    Gaussian: $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$;
  2. 2.
    asymmetric Laplace (ALD) with parameter $\tau=0.5$, which aligns the conditional median exactly with $Q_{\tau}(\cdot)$;
  3. 3.
    Student-$t_{3}$ noise;
  4. 4.
    contaminated normal: a $90\%$$10\%$ mixture of $\mathcal{N}(0,\sigma^{2})$ and $\mathcal{N}(0,25\sigma^{2})$;
  5. 5.
    Cauchy errors;
  6. 6.
    heteroskedastic $t_{3}$ errors,
    $\varepsilon_{i}=\sigma(u_{i})\eta_{i},\qquad\sigma(u_{i})=0.5+0.5u_{i1},\ \eta_{i}\sim t_{3},$
    which induce a left–right gradient in noise scale.
For the contaminated normal and Cauchy cases we additionally flag as outliers those observations with $|\varepsilon_{i}|$ exceeding a high cut-off (e.g. $2.5$ times the Gaussian standard deviation). These outliers are indicated as hollow circles overlaid on the true fields in the plots, illustrating that the proposed quantile-based estimator is robust to a small number of extreme data points.
For each error scenario, we fit SS–SVCQR to $\tau=0.5$ using the spatially blocked cross-validation scheme of Section 3 to select $(\lambda_{1},\lambda_{2})$, and then compare the estimated deviations $\hat{\delta}_{j}(u_{i})$ with the truth.
Figures 16 show, for each covariate $X_{j}$, three panels: the true deviation $\delta_{0,j}(u_{i})$, the estimated deviation $\hat{\delta}_{j}(u_{i})$, and their difference (error). All panels share a common color scale so that amplitude and sign can be compared directly across variables and scenarios. In all cases $X_{2}$ and $X_{4}$ appear essentially flat, confirming that the method does not spuriously create spatial variation for truly global covariates, while the complex patterns for $X_{1}$ and $X_{3}$ are well recovered even under heavy-tailed and heteroskedastic noise.
Figure 1: Scenario 1 under normal errors.
Figure 2: Scenario 1 under asymmetric Laplace errors.
Figure 3: Scenario 1 under Student-$t_{3}$ errors.
Figure 4: Scenario 1 under contaminated normal errors (hollow circles mark gross outliers).
Figure 5: Scenario 1 under Cauchy errors (hollow circles mark extreme outliers).
Figure 6: Scenario 1 under heteroskedastic $t_{3}$ errors with spatially varying scale.
Figures 16 summarize the results in the six error scenarios. In each figure, the rows correspond to covariates $X_{1},\dots,X_{4}$, while the four columns, from left to right, show the true spatial deviation, the estimated deviation from SS–SVCQR, the pointwise error and the total effect $\beta_{G,j}+\delta_{j}(u)$ at the target quantile.
The first column shows the simulated ground truth: the deviation field $\delta_{0,j}(u)$ evaluated at the sampled locations, together with smoothed contour lines that make the underlying spatial pattern easier to discern. For $X_{1}$ this reveals an oscillatory wave-like pattern superimposed on a mild trend, while $X_{3}$ exhibits a smooth dome-shaped hotspot; by contrast, $X_{2}$ and $X_{4}$ are truly global and their deviation fields are essentially flat, apart from the overlaid outlier points in the heavy-tailed scenarios.
The second column contains the corresponding estimates $\hat{\delta}_{j}(u)$ and shows that the proposed estimator reproduces the main qualitative features of the true fields: the sign changes, level sets and apparent boundaries for $X_{1}$ and $X_{3}$ closely track those in the first column, while the panels for $X_{2}$ and $X_{4}$ remain visually featureless, indicating that spurious spatial variation is not introduced for covariates that are in fact global.
The third column displays the difference $\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})$; here the colors fluctuate around zero without forming coherent spatial patches, and there is no obvious residual spatial correlation. This noise-like appearance of the error field indicates that the dominant spatial structure has been absorbed by the model and that the remaining lack-of-fit behaves as idiosyncratic noise rather than as a missing large-scale trend.
Finally, the fourth column reports the total effect of each covariate in location $u_{i}$, combining the global component and the spatial deviation. For $X_{1}$ and $X_{3}$ the panels are dominated by warm colors, suggesting an overall positive association if only a single global coefficient were to be examined; however, the first column reveals sizeable regions where the local deviation is negative and partially cancels the global effect. Without explicitly separating a global level from a spatial deviation, a purely global quantile regression would report a single positive coefficient and completely miss these local reversals, potentially leading to misleading substantive conclusions.
Across all six error mechanisms, including heavy-tailed, contaminated, and heteroskedastic designs, the same qualitative picture emerges: the estimated deviation fields in the second column closely match the truth, the error fields in the third column are spatially structureless, and the total-effect panels in the fourth column make clear that spatial nonstationarity can be substantial even when the overall effect is positive on average.
5.2 Monte Carlo Assessment
Scenario 2 quantifies performance over repeated samples. The spatial design, covariates and coefficient surfaces are exactly as in the common setup above, but we now generate $R=100$ independent datasets with sample size $n=1000$ for each error scenario of interest (e.g. Gaussian, ALD, Student-$t_{3}$, and Cauchy). For each replicate we fit SS–SVCQR at $\tau=0.5$ with cross-validated $(\lambda_{1},\lambda_{2})$ and record:
  • 1.
    Global parameter error
    $\mathrm{PE}_{\theta}=\|\hat{\bm{\alpha}}-\bm{\alpha}_{0}\|_{2}+\|\hat{\bm{\beta}}_{G}-\bm{\beta}_{G0}\|_{2}.$
  • 2.
    Deviation mean-squared error for each covariate,
    $\mathrm{MSE}_{\delta,j}=\frac{1}{n}\sum_{i=1}^{n}\big\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\big\}^{2},\qquad j=1,\dots,4.$
  • 3.
    Local/global classification accuracy. We declare $X_{j}$ to be “local” if
    $\sqrt{\frac{1}{n}\sum_{i=1}^{n}\hat{\delta}_{j}(u_{i})^{2}}>\kappa$
    for a small threshold $\kappa$, and “global” otherwise. Using the true set of local covariates $\{X_{1},X_{3}\}$, we compute the number of true positives (TP), true negatives (TN), false positives (FP) and false negatives (FN), together with sensitivity $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$ and specificity $\mathrm{TN}/(\mathrm{TN}+\mathrm{FP})$.
  • 4.
    Predictive performance measured by the average check loss on an independent test set of size $n_{\mathrm{test}}$,
    $\mathrm{CL}_{\mathrm{test}}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\rho_{\tau}\Big(Y_{i}^{\mathrm{test}}-\hat{Q}_{\tau}(Y_{i}^{\mathrm{test}}\mid\bm{Z}_{i}^{\mathrm{test}},\bm{X}_{i}^{\mathrm{test}},u_{i}^{\mathrm{test}})\Big).$
In order to objectively evaluate the performance of the proposed SS–SVCQR model, we compare it with five representative competitors. (i) Global quantile regression (QR) uses a purely global specification and thus ignores spatial nonstationarity. (ii) Standard geographically weighted regression (GWR) treats all covariates as spatially varying. (iii) Mixed GWR assumes that the local and global covariates are known a priori and fit a partially varying model; this can be regarded as an oracle benchmark for methods that must learn the local/global structure. (iv) Quantile generalized additive models (QGAM) allow for flexible nonlinear effects but do not directly target spatially varying coefficients. (v) The classical spatially varying coefficient (SVC) model is fitted by least squares without quantile weighting. These methods span the spectrum from purely global to fully spatially varying specifications and from mean to quantile regression, and they are chosen to assess three aspects: (a) accuracy of global and varying coefficient estimation, (b) ability to recover the true local versus global covariates, and (c) predictive performance under a range of error distributions. We report the detailed results of the Monte Carlo simulation in the six tables below. [^1] In Tables 16, the values in parentheses represent the standard deviation corresponding to that metric.
Table 1: Monte Carlo result for quantile regression (QR)
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.571 (0.157) 6.824 (0.359) 0.021 (0.026) 3.081 (0.166) 0.024 (0.036) 0 (0) 1 (0) 3.040 (0.337)
ALD 0.740 (0.215) 6.837 (0.350) 0.039 (0.060) 3.099 (0.188) 0.043 (0.062) 0 (0) 1 (0) 3.039 (0.364)
hetero 0.431 (0.117) 1.755 (0.244) 0.080 (0.027) 0.631 (0.117) 0.077 (0.028) 1 (0) 0 (0) 2.847 (0.406)
t3 0.587 (0.162) 6.814 (0.363) 0.019 (0.030) 3.089 (0.183) 0.024 (0.032) 0 (0) 1 (0) 2.998 (0.370)
contam 0.590 (0.188) 6.824 (0.368) 0.020 (0.034) 3.085 (0.177) 0.023 (0.026) 0 (0) 1 (0) 3.005 (0.317)
cauchy 0.701 (0.215) 6.839 (0.361) 0.033 (0.038) 3.100 (0.180) 0.032 (0.049) 0 (0) 1 (0) 5.803 (38.59)
Table 2: Standard_GWR
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.471 (0.108) 2.155 (0.207) 0.064 (0.031) 0.705 (0.135) 0.065 (0.028) 1 (0) 0 (0) 2.844 (0.342)
ALD 0.615 (0.186) 2.233 (0.244) 0.152 (0.071) 0.799 (0.196) 0.134 (0.065) 1 (0) 0 (0) 2.956 (0.401)
hetero 0.483 (0.115) 2.173 (0.206) 0.071 (0.040) 0.728 (0.129) 0.070 (0.039) 1 (0) 0 (0) 2.809 (0.405)
t3 0.508 (0.122) 2.189 (0.216) 0.083 (0.046) 0.743 (0.135) 0.083 (0.044) 1 (0) 0 (0) 2.825 (0.403)
contam 0.516 (0.129) 2.179 (0.230) 0.090 (0.042) 0.730 (0.144) 0.090 (0.037) 1 (0) 0 (0) 2.877 (0.326)
cauchy 124.1 (1099.3) 9.3e59.3105 (9.3e69.3106) 1.2e51.2105 (1.2e61.2106) 8.9e58.9105 (8.9e68.9106) 1.5e51.5105 (1.5e61.5106) 1 (0) 0 (0) 40.78 (305.3)
Table 3: Mixed_GWR
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.255 (0.063) 0.568 (0.117) 0.094 (0.022) 0.238 (0.053) 0.096 (0.027) 1 (0) 0 (0) 2.869 (0.384)
ALD 0.505 (0.159) 1.025 (0.251) 0.259 (0.087) 0.500 (0.154) 0.236 (0.087) 1 (0) 0 (0) 2.994 (0.404)
hetero 0.284 (0.087) 0.637 (0.147) 0.112 (0.038) 0.277 (0.075) 0.113 (0.044) 1 (0) 0 (0) 2.817 (0.411)
t3 0.337 (0.102) 0.723 (0.166) 0.142 (0.050) 0.322 (0.085) 0.146 (0.063) 1 (0) 0 (0) 2.853 (0.391)
contam 0.353 (0.098) 0.753 (0.174) 0.156 (0.056) 0.328 (0.076) 0.159 (0.044) 1 (0) 0 (0) 2.901 (0.361)
cauchy 119.5 (1062.8) 2.0e52.0105 (2.0e62.0106) 2.8e42.8104 (2.7e52.7105) 2.1e52.1105 (2.1e62.1106) 3.3e43.3104 (3.3e53.3105) 1 (0) 0 (0) 42.63 (333.9)
Table 4: QGAM
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.248 (0.060) 1.480 (0.152) 0.019 (0.015) 0.121 (0.036) 0.020 (0.016) 1 (0) 0 (0) 2.992 (0.381)
ALD 0.427 (0.140) 1.582 (0.182) 0.062 (0.048) 0.246 (0.094) 0.060 (0.050) 1 (0) 0 (0) 3.018 (0.428)
hetero 0.230 (0.068) 1.481 (0.147) 0.016 (0.015) 0.118 (0.035) 0.017 (0.016) 1 (0) 0 (0) 2.971 (0.420)
t3 0.270 (0.076) 1.484 (0.144) 0.022 (0.019) 0.134 (0.044) 0.023 (0.020) 1 (0) 0 (0) 2.969 (0.405)
contam 0.257 (0.063) 1.508 (0.156) 0.022 (0.017) 0.129 (0.035) 0.021 (0.019) 1 (0) 0 (0) 3.001 (0.352)
cauchy 0.578 (1.299) 3.076 (11.19) 0.229 (1.675) 1.309 (9.401) 0.317 (2.566) 1 (0) 0 (0) 5.814 (38.61)
Table 5: Classic_SVC
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.272 (0.066) 1.454 (0.142) 0.025 (0.020) 0.129 (0.042) 0.026 (0.020) 1 (0) 0 (0) 2.989 (0.377)
ALD 0.467 (0.151) 1.595 (0.193) 0.077 (0.052) 0.274 (0.108) 0.075 (0.056) 1 (0) 0 (0) 3.026 (0.421)
t3 0.316 (0.091) 1.488 (0.145) 0.035 (0.035) 0.171 (0.098) 0.041 (0.049) 1 (0) 0 (0) 2.962 (0.410)
contam 0.333 (0.088) 1.502 (0.146) 0.039 (0.036) 0.164 (0.048) 0.036 (0.028) 1 (0) 0 (0) 2.987 (0.365)
cauchy 114.8 (1016.2) 4.9e54.9105 (4.9e64.9106) 7.0e47.0104 (7.0e57.0105) 5.0e55.0105 (5.0e65.0106) 8.4e48.4104 (8.4e58.4105) 1 (0) 0 (0) 47.69 (371.2)
Table 6: $S$-SVCQR
MSE
Error PEθ $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ Sens. Spec. CL
normal 0.098 (0.027) 0.195 (0.014) 0 (0) 0.159 (0.010) 0 (0) 1 (0) 1 (0) 0.474 (0.009)
ALD 0.193 (0.054) 0.405 (0.035) 0 (0) 0.341 (0.038) 0 (0) 1 (0) 0.985 (0.086) 1.075 (0.023)
hetero 0.086 (0.021) 0.182 (0.012) 0 (0) 0.150 (0.009) 0 (0) 1 (0) 1 (0) 0.494 (0.012)
t3 0.113 (0.028) 0.234 (0.016) 0 (0) 0.194 (0.012) 0 (0) 1 (0) 1 (0) 0.630 (0.015)
contam 0.107 (0.032) 0.226 (0.016) 0 (0) 0.186 (0.015) 0 (0) 1 (0) 1 (0) 0.634 (0.018)
cauchy 0.156 (0.041) 0.333 (0.030) 0 (0) 0.277 (0.030) 0 (0) 1 (0) 0.990 (0.070) 6.611 (21.23)
Tables 16 summarise Monte Carlo means and standard deviations of the above metrics over $R=500$ replicates in Scenario 2. For the purely global quantile regression (QR) in Table 1, the global parameter error $\mathrm{PE}_{\theta}$ and deviation MSEs are substantially larger than those of the spatial competitors whenever spatial nonstationarity is present. Moreover, QR classifies all covariates as global, leading to sensitivity essentially equal to zero and specificity equal to one, and its predictive performance as measured by $\mathrm{CL}_{\mathrm{test}}$ is clearly inferior once the error distribution departs from the Gaussian case.
Standard GWR and mixed GWR (Tables 2 and 3) reduce $\mathrm{PE}_{\theta}$ and the deviation MSEs relative to QR, but they do so by treating almost all effects as spatially varying. As a consequence, sensitivity is essentially one while specificity is close to zero across all error distributions, indicating severe overfitting of local structure. QGAM (Table 4) shows a similar pattern: it achieves smaller $\mathrm{PE}_{\theta}$ and deviation MSEs than QR, but again classifies nearly all covariates as local, with sensitivity near one and specificity near zero. The classical SVC model in Table 5 behaves analogously, and under heavy-tailed Cauchy errors its parameter error and deviation MSEs can become extremely large despite reasonably small check loss in some scenarios.
In contrast, the proposed SS–SVCQR estimator in Table 6 delivers uniformly small global parameter errors and deviation MSEs across all error settings. With the larger penalty used here, SS–SVCQR achieves sensitivity essentially equal to one and specificity close to one for normal, heteroscedastic, $t_{3}$, and contaminated-normal errors, and only slightly below one for ALD and Cauchy errors. Thus SS–SVCQR is the only method that simultaneously recovers both truly local covariates $\{X_{1},X_{3}\}$ and truly global covariates $\{X_{2},X_{4}\}$ with high accuracy. Its predictive check loss $\mathrm{CL}_{\mathrm{test}}$ is uniformly lower than, or comparable to, that of the best competitor for light- and moderately heavy-tailed errors, and remains stable even under Cauchy errors, where several fully spatial methods suffer from dramatic loss of efficiency. These results confirm that SS–SVCQR provides near-oracle estimation of global coefficients, accurate recovery of local structure, and competitive predictive performance over a wide range of error distributions in Scenario 2.
Monte Carlo averages and standard deviations of these metrics show that SS–SVCQR has high sensitivity and specificity for distinguishing local (varying) from global (constant) covariates, even under heavy-tailed and heteroskedastic errors; the deviation MSEs remain small relative to the signal variance, and the global coefficients are estimated with near-oracle precision. Predictive performance is competitive with, or better than, fully spatially varying quantile regression and markedly improves over a purely global quantile regression when spatial nonstationarity is present.
6 Real Data
Appendix A Proofs of Theorems
Throughout the appendix we keep the notation of Sections 24. For a vector $\bm{v}\in\mathbb{R}^{n}$, write $\|\bm{v}\|_{2}^{2}=n^{-1}\sum_{i=1}^{n}v_{i}^{2}\cdot n$ when we need to emphasize averaging and use $C,c,c_{0},c_{1},\dots$ for generic positive constants, possibly changing from line to line. We denote by $\langle\bm{a},\bm{b}\rangle=\bm{a}^{\top}\bm{b}$ the Euclidean inner product and by $\preceq$ the Löwner order for symmetric matrices. Recall the diagonal multipliers $\bm{X}_{\odot j}=\operatorname{diag}(X_{1j},\dots,X_{nj})$, the Laplacian $\bm{L}_{\mathrm{sym}}$ in (4), the degree matrix $\bm{D}$, and the degree–weighted centering operator $\mathsf{Proj}_{D}$ in (10). The subgradient of the check loss is $\psi_{\tau}(r)=\tau-\mathbf{1}\{r<0\}$ and we use $\bm{\psi}_{\tau}(\bm{r})=(\psi_{\tau}(r_{1}),\dots,\psi_{\tau}(r_{n}))^{\top}$.
We repeatedly use two standard tools.
(i) Knight identity and local quadratic expansion.
For any $r,t\in\mathbb{R}$,
$\rho_{\tau}(r-t)-\rho_{\tau}(r)=-t\psi_{\tau}(r)+\int_{0}^{t}\big\{\mathbf{1}(r\leq s)-\mathbf{1}(r\leq 0)\big\}ds.$ (15)
Under Assumption 3, taking conditional expectations and using a Taylor expansion of the conditional cdf at 0 yields, uniformly for $\|t\|_{\infty}=o(1)$,
$\mathrm{E}\left\{\rho_{\tau}(\varepsilon_{i}-t)-\rho_{\tau}(\varepsilon_{i})\mid\bm{Z}_{i},\bm{X}_{i},u_{i}\right\}=\frac{1}{2}f_{\varepsilon\mid Z,X,U}(0)t^{2}+o(t^{2}).$ (16)
See, e.g., [Koenker2005] and the original derivation in [Knight1998].
(ii) Graph Poincaré inequality on the centered subspace.
Let $\mathcal{S}=\{\bm{v}\in\mathbb{R}^{n}:\ \bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{v}=0\ \text{for all components }\mathcal{C}\}$. Then $\mathcal{S}$ is orthogonal (in the $\bm{D}$–inner product) to the null space of $\bm{L}_{\mathrm{sym}}$. Since $\bm{L}_{\mathrm{sym}}$ is symmetric p.s.d. with eigenvalues in $[0,2]$, there exists $c_{\lambda}>0$ (depending on the graph but not on $n$ under Assumption 1) such that
$\bm{v}^{\top}\bm{L}_{\mathrm{sym}}\bm{v}\ \geq\ c_{\lambda}\|\bm{v}\|_{2}^{2}\quad\text{for all }\ \bm{v}\in\mathcal{S}.$ (17)
See [Chung1997, Chap. 1] and [vonLuxburg2007, Sec. 2].
Let $(\hat{\bm{\theta}}_{P},\hat{\bm{\Theta}}_{L})$ be any global minimizer of $\mathcal{L}_{n}$ in Section 4. For any feasible $(\bm{\theta}_{P},\bm{\Theta}_{L})$ we have the basic inequality
$0$ $\leq\mathcal{L}_{n}(\bm{\theta}_{P},\bm{\Theta}_{L})-\mathcal{L}_{n}(\hat{\bm{\theta}}_{P},\hat{\bm{\Theta}}_{L})$
$=\sum_{i=1}^{n}\Big\{\rho_{\tau}\big(Y_{i}-q_{\tau,i}(\bm{\theta}_{P},\bm{\Theta}_{L})\big)-\rho_{\tau}\big(Y_{i}-q_{\tau,i}(\hat{\bm{\theta}}_{P},\hat{\bm{\Theta}}_{L})\big)\Big\}$
$\quad+\sum_{j=1}^{p}\lambda_{1n}w_{j,n}\big(\|\bm{\delta}_{j}\|_{2}-\|\hat{\bm{\delta}}_{j}\|_{2}\big)+\sum_{j=1}^{p}\lambda_{2n}\big(\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j}-\hat{\bm{\delta}}_{j}^{\top}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}\big),$ (18)
where $q_{\tau,i}(\cdot)$ is the $i$th component of the conditional quantile vector in (7). Choosing $(\bm{\theta}_{P},\bm{\Theta}_{L})=(\bm{\theta}_{P0},\bm{\Theta}_{L0})$ and expanding the check loss via (15) around the true residuals $\varepsilon_{i}$ give the starting point for the deviation analysis.
Proof of Theorem 3.
Write $\Delta\bm{\alpha}=\hat{\bm{\alpha}}-\bm{\alpha}_{0}$, $\Delta\bm{\beta}_{G}=\hat{\bm{\beta}}_{G}-\bm{\beta}_{G0}$ and $\Delta\bm{\delta}_{j}=\hat{\bm{\delta}}_{j}-\bm{\delta}_{j0}$, and define the local perturbation $\bm{h}=\sum_{j=1}^{p}\bm{X}_{\odot j}\Delta\bm{\delta}_{j}$. By feasibility, each $\Delta\bm{\delta}_{j}\in\mathcal{S}$ (componentwise degree–weighted centering). Let $\bm{r}_{0}=\bm{\varepsilon}$ be the true residual vector. Using (15) with $t_{i}=h_{i}+\bm{Z}_{i}^{\top}\Delta\bm{\alpha}+\bm{X}_{i}^{\top}\Delta\bm{\beta}_{G}$ and summing, we obtain
$\sum_{i=1}^{n}\Big\{\rho_{\tau}(\varepsilon_{i}-t_{i})-\rho_{\tau}(\varepsilon_{i})\Big\}=-\langle\bm{\psi}_{\tau}(\bm{\varepsilon}),\bm{t}\rangle+\sum_{i=1}^{n}\int_{0}^{t_{i}}\big\{\mathbf{1}(\varepsilon_{i}\leq s)-\mathbf{1}(\varepsilon_{i}\leq 0)\big\}ds.$ (19)
By Assumption 7 (local quadratic expansion) and bounded covariates (Assumption 2), uniformly on events where $\|\Delta\bm{\theta}_{P}\|_{2}=O_{p}(n^{-1/2})$,
$\sum_{i=1}^{n}\Big\{\rho_{\tau}(\varepsilon_{i}-t_{i})-\rho_{\tau}(\varepsilon_{i})\Big\}=\tfrac{1}{2}\bm{t}^{\top}\bm{W}\bm{t}-\langle\bm{\psi}_{\tau}(\bm{\varepsilon}),\bm{t}\rangle+o_{p}(\|\bm{t}\|_{2}^{2}),$ (20)
where $\bm{W}=\operatorname{diag}\{f_{\varepsilon\mid Z,X,U}(0),\dots\}$ has entries bounded between $c_{f}$ and $C_{f}$.
Part (i): mean-squared rate.
Insert $(\bm{\theta}_{P},\bm{\Theta}_{L})=(\bm{\theta}_{P0},\bm{\Theta}_{L0})$ into (18). Combine (20) with polarization for the Laplacian term to obtain for every $j$,
$\tfrac{1}{2}\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{W}^{2}+\lambda_{2n}\Big(\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}-\hat{\bm{\delta}}_{j}^{\top}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}\Big)+\lambda_{1n}w_{j,n}\big(\|\bm{\delta}_{j0}\|_{2}-\|\hat{\bm{\delta}}_{j}\|_{2}\big)$
$\qquad\leq\langle\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon}),\Delta\bm{\delta}_{j}\rangle+R_{n,j},$ (21)
with $R_{n,j}=o_{p}(\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2})$ collecting the parametric perturbations and higher-order terms. Using
$\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}-\hat{\bm{\delta}}_{j}^{\top}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}=-\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\Delta\bm{\delta}_{j}-2\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0},$
the centering $\Delta\bm{\delta}_{j}\in\mathcal{S}$ and (17) yield
$\tfrac{1}{2}c_{f}\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}+\lambda_{2n}c_{\lambda}\|\Delta\bm{\delta}_{j}\|_{2}^{2}\ \leq\ \langle\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon}),\Delta\bm{\delta}_{j}\rangle+2\lambda_{2n}\big|\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big|+o_{p}(\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}).$ (22)
By Cauchy–Schwarz and Lemma-type bounds for $\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\bm{\varepsilon})$ (bounded covariates, Assumption 2), the stochastic term is $O_{p}(\sqrt{n}\|\Delta\bm{\delta}_{j}\|_{2})$. The bias term satisfies
$\big|\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big|\leq\big(\Delta\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\Delta\bm{\delta}_{j}\big)^{1/2}\big(\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}\big)^{1/2}\leq C\|\Delta\bm{\delta}_{j}\|_{2}\mathcal{S}_{j}^{1/2},$
with $\mathcal{S}_{j}=\bm{\delta}_{j0}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j0}$ finite by Assumption 4. Absorbing the $o_{p}(\cdot)$ and dividing by $\|\Delta\bm{\delta}_{j}\|_{2}$,
$\lambda_{2n}c_{\lambda}\|\Delta\bm{\delta}_{j}\|_{2}\ \leq\ C\Big(\sqrt{n}+\lambda_{2n}\mathcal{S}_{j}^{1/2}\Big),$
so that
$\|\Delta\bm{\delta}_{j}\|_{2}^{2}\ \leq\ C\left(\frac{n}{\lambda_{2n}^{2}}+\mathcal{S}_{j}\right).$
Finally, using boundedness of $X_{ij}$ (Assumption 2) to relate $\|\bm{X}_{\odot j}\Delta\bm{\delta}_{j}\|_{2}^{2}$ and $\|\Delta\bm{\delta}_{j}\|_{2}^{2}$, we obtain
$\frac{1}{n}\sum_{i=1}^{n}\big\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\big\}^{2}=O_{p}\Big(\frac{1}{n\lambda_{2n}}+\lambda_{2n}\mathcal{S}_{j}\Big),$
and the bound is balanced by $\lambda_{2n}\asymp(n\mathcal{S}_{j})^{-1/2}$, proving (i).
Part (ii): group selection consistency.
Consider the KKT conditions for each $\hat{\bm{\delta}}_{j}$:
$\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})+2\lambda_{2n}\bm{L}_{\mathrm{sym}}\hat{\bm{\delta}}_{j}+\bm{D}\hat{\bm{\mu}}_{j}\in-\lambda_{1n}w_{j,n}\partial\|\hat{\bm{\delta}}_{j}\|_{2},\qquad\bm{1}_{\mathcal{C}}^{\top}\bm{D}\hat{\bm{\delta}}_{j}=0.$
Inactive groups. Let $j\in\mathcal{A}_{0}$. Test $\hat{\bm{\delta}}_{j}=\bm{0}$. Projecting onto the centered subspace (which eliminates $\bm{D}\hat{\bm{\mu}}_{j}$) and using $\|\mathsf{Proj}_{D}(\bm{v})\|_{2}\leq\|\bm{v}\|_{2}$ yields
$\Big\|\mathsf{Proj}_{D}\big(\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\big)\Big\|_{2}\ \leq\ \|\bm{X}_{\odot j}^{\top}\bm{\psi}_{\tau}(\hat{\bm{r}})\|_{2}=O_{p}(\sqrt{n})\ \leq\ \lambda_{1n}w_{j,n}$
with probability tending to one because $\sqrt{n}\lambda_{1n}\to\infty$ and $\inf_{j\in\mathcal{A}_{0}}w_{j,n}\geq c_{w}>0$ by Assumption 5. Hence $\Pr(\hat{\bm{\delta}}_{j}=\bm{0})\to 1$ for each $j\in\mathcal{A}_{0}$.
Active groups. Let $j\in\mathcal{A}_{0}^{c}$. Consider the restricted problem that sets $\bm{\delta}_{j}=\bm{0}$ for $j\in\mathcal{A}_{0}$ and optimizes over $\{\bm{\delta}_{j}:j\in\mathcal{A}_{0}^{c}\}$ and $\bm{\theta}_{P}$. Under Assumptions 56, the groupwise beta–min separation and bounded adaptive weights on $\mathcal{A}_{0}^{c}$ ensure that the restricted solution satisfies $\|\hat{\bm{\delta}}_{j}^{\mathrm{res}}\|_{2}\geq c_{\beta}\lambda_{1n}w_{j,n}-o_{p}(\lambda_{1n})>0$ for $j\in\mathcal{A}_{0}^{c}$, while the subgradient conditions for $j\in\mathcal{A}_{0}$ hold with strict inequality (dual feasibility). Therefore the restricted solution is also optimal for the full problem (primal–dual witness, cf. [YuanLin2006, Simon2013SGL]), implying $\hat{\bm{\delta}}_{j}\neq\bm{0}$ for all $j\in\mathcal{A}_{0}^{c}$. On this event, the rate in part (i) holds for $j\in\mathcal{A}_{0}^{c}$. ∎
Proof of Lemma 1.
Let $e_{j,i}=\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})$ and write $\bm{e}_{j}=\hat{\bm{\delta}}_{j}-\bm{\delta}_{j0}$. Using Cauchy–Schwarz, boundedness of $f_{\varepsilon\mid Z,X,U}(0)$ and of covariates (Assumption 2), and denoting $\bm{G}_{i}=(\bm{Z}_{i}^{\top},\bm{X}_{i}^{\top})^{\top}$,
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}\right\|_{2}$ $\leq C\left(\frac{1}{n}\sum_{i=1}^{n}\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)^{2}\right)^{1/2}\left(\frac{1}{n}\sum_{i=1}^{n}\|\bm{G}_{i}\|_{2}^{2}\right)^{1/2}$
$\leq C^{\prime}\left(\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}e_{j,i}^{2}\right)^{1/2}=C^{\prime}\left(\frac{1}{p}\sum_{j=1}^{p}\frac{1}{n}\sum_{i=1}^{n}e_{j,i}^{2}\right)^{1/2}.$
By Theorem 3(i), for each $j$,
$\frac{1}{n}\sum_{i=1}^{n}e_{j,i}^{2}=O_{p}\Big(\frac{1}{n\lambda_{2n}}+\lambda_{2n}\mathcal{S}_{j}\Big).$
Averaging over $j$ preserves the order (with $\bar{\mathcal{S}}=p^{-1}\sum_{j}\mathcal{S}_{j}$), hence
$\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}\right\|_{2}=O_{p}\left(\Big(\tfrac{1}{n\lambda_{2n}}+\lambda_{2n}\bar{\mathcal{S}}\Big)^{1/2}\right).$
For the canonical balancing $\lambda_{2n}\asymp n^{-1/2}$, the right-hand side is $O_{p}(n^{-1/4})=o_{p}(1)$, which proves the claim. ∎
Proof of Theorem 4.
Let $\hat{\bm{r}}$ be the fitted residuals and $r_{0,i}=\varepsilon_{i}$. The stationarity of the parametric block gives
$\bm{0}=\sum_{i=1}^{n}\psi_{\tau}(\hat{r}_{i})\bm{G}_{i}=\sum_{i=1}^{n}\psi_{\tau}(r_{0,i})\bm{G}_{i}+\sum_{i=1}^{n}\big\{\psi_{\tau}(\hat{r}_{i})-\psi_{\tau}(r_{0,i})\big\}\bm{G}_{i}.$
Using the linearization
$\psi_{\tau}(\varepsilon_{i}-t_{i})-\psi_{\tau}(\varepsilon_{i})=-f_{\varepsilon\mid Z,X,U}(0)t_{i}+\eta_{i}(t_{i}),$
with $\sum_{i}\eta_{i}(t_{i})\bm{G}_{i}=o_{p}(\sqrt{n}\|\bm{t}\|_{2})$ uniformly for $\|\bm{t}\|_{2}=O_{p}(1)$ (see [Koenker2005, Sec. 4.3]), we obtain
$\bm{0}$ $=\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}-\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)t_{i}\bm{G}_{i}+o_{p}(\sqrt{n}),$ (23)
$t_{i}$ $=\bm{Z}_{i}^{\top}\Delta\bm{\alpha}+\bm{X}_{i}^{\top}\Delta\bm{\beta}_{G}+\sum_{j=1}^{p}X_{ij}\big\{\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})\big\}.$ (24)
Divide (23) by $\sqrt{n}$ and rearrange the terms that involve $(\Delta\bm{\alpha},\Delta\bm{\beta}_{G})$:
$\sqrt{n}\bm{M}\left[\begin{array}[]{c}\Delta\bm{\alpha}\\ \Delta\bm{\beta}_{G}\end{array}\right]=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f_{\varepsilon\mid Z,X,U}(0)\Big(\sum_{j=1}^{p}X_{ij}e_{j,i}\Big)\bm{G}_{i}+o_{p}(1),$
where $\bm{M}=\mathrm{E}[f_{\varepsilon\mid Z,X,U}(0)\bm{G}_{i}\bm{G}_{i}^{\top}]$ (Assumption 7) and $e_{j,i}=\hat{\delta}_{j}(u_{i})-\delta_{0,j}(u_{i})$. By Lemma 1, the middle term on the right is $o_{p}(1)$ after multiplication by $\bm{M}^{-1}$. Therefore
$\sqrt{n}\left[\begin{array}[]{c}\Delta\bm{\alpha}\\ \Delta\bm{\beta}_{G}\end{array}\right]=\bm{M}^{-1}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\tau}(\varepsilon_{i})\bm{G}_{i}\right)+o_{p}(1).$
By a multivariate central limit theorem, the leading sum converges in distribution to $\mathcal{N}(\bm{0},\bm{V})$ with $\bm{V}=\tau(1-\tau)\mathrm{E}(\bm{G}_{i}\bm{G}_{i}^{\top})$, yielding
$\sqrt{n}(\hat{\bm{\theta}}_{P}-\bm{\theta}_{P0})\ \overset{d}{\longrightarrow}\ \mathcal{N}(\bm{0},\ \bm{M}^{-1}\bm{V}\bm{M}^{-1}).$
Oracle property. By Theorem 3(ii), $\Pr(\hat{\mathcal{A}}=\mathcal{A}_{0})\to 1$. On this event, $\hat{\bm{\delta}}_{j}=\bm{0}$ for $j\in\mathcal{A}_{0}$ and the restricted problem for $(\bm{\theta}_{P},\{\bm{\delta}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ coincides with the oracle problem that knows $\mathcal{A}_{0}$. Hence the joint limit of $(\hat{\bm{\theta}}_{P},\{\hat{\bm{\delta}}_{j}\}_{j\in\mathcal{A}_{0}^{c}})$ equals the oracle limit distribution. ∎